# Distinguishing Power of 4-Legendrian Permutation Racks

### v1

@@ 5.1 Legendrian Knots
Legendrian knots are important objects of study in the field of contact topology. We briefly review several concepts from the theory; we refer the reader to the survey of Etnyre [4] for a rigorous treatment.
The standard contact structure is the kernel of the differential 1-form dzydxdz-y\,dx in $\mathbb{R}^{3}$, which is depicted in Figure 1. A smooth knot in $\mathbb{R}^{3}$ is called Legendrian if it lies everywhere tangent to the standard contact structure.
@@
Legendrian knots are usually studied via their front projections to the $xz$-plane. Front projections of Legendrian knots have two key features that distinguish them from projections of smooth knots. First, since tangent lines can never be vertical, front projections have cusps in place of vertical tangencies. Second, due to the direction of twisting, the overstrand at each crossing is always the strand having the more negative slope. Note that an oriented front projection never has two leftward-oriented or rightward-oriented cusps placed adjacent to each other. For example, Figure 3 depicts a Legendrian left handed trefoil in its front projection.
@@
However, there exist many nonequivalent Legendrian knots that share the same topological knot type and classical invariants. This motivates the search for Legendrian knot invariants that distinguish such pairs of Legendrian knots.
@@
In particular, Kimura [10] and Cheng and He [2] posed the open questions of whether GL-racks and 4-Legendrian racks provide effective invariants of Legendrian knots. While this question remains open in general, Kimura [11, Thm. 4.2.3] gave a negative answer for all 4-Legendrian quandles, and our main theorem (Theorem 6.3) provides a negative answer for all 4-Legendrian permutation racks.
@@ 5.2 Coloring Invariants
In [11, Sec. 4.2], Kimura introduced an invariant of Legendrian knots $L$ called the fundamental 4-Legendrian rack, which we denote by $F(L)$. This is defined in the same vein as the fundamental GL-rack, which Karmakar, Saraf, and Singh introduced in [9, Thm. 4.3], and the fundamental quandle of a smooth knot, which was independently introduced by Joyce [8] and Matveev [13].
We outline the construction here; see [11, Sec. 4.2] for details, and cf. [9, 2]. 4-Legendrian racks can be viewed as an algebraic theory with two binary operations $\triangleright^{\pm 1}$ and four unary operations $u_{l}^{\pm 1},u_{r}^{\pm 1}$. It follows from general results from universal algebra that we can consider free 4-Legendrian racks, which are defined using the usual universal property, and take quotients of 4-Legendrian racks by congruence relations.
Fix an oriented front projection of a Legendrian knot $L$. Label the arcs (i.e., connected components) of the front projection by $x_{1},\dots,x_{n}$, and let $F\coloneq\langle x_{1},\dots,x_{n}\rangle$ be the free 4-Legendrian rack generated by $x_{1},\dots,x_{n}$. Starting at any crossing in the front projection, traverse the knot using its given orientation. At each cusp, impose a label $u_{l},u_{r},d_{l}$, or $d_{r}$ in the obvious way, depending on the orientation of the cusp; cf. (4.2). At each crossing, impose a relation on $F$ between arcs as illustrated in Figure 2. Define $F(L)$ to be $F$ modulo the congruence relation generated by these $n$ crossing relations.
@@
It can be shown that the isomorphism class of $F(L)$ is an invariant of $L$; see [11, Sec. 4.2]. Of course, this invariant is difficult to use directly, so we instead consider colorings of $F(L)$ by finite 4-Legendrian racks. These are easier to compute since $F(L)$ is finitely presented (having $n$ generators and $n$ relations in the above construction).
Definition 5.4.
@@
Since $F(L)$ is an invariant of $L$, the set $\operatorname{Hom}(F(L),X)$ of colorings of $L$ by a given 4-Legendrian rack $(X,\triangleright,u_{l},u_{r})$ is also an invariant of $L$; see [11, Rem. 30].
@@
Consider the Legendrian left handed trefoil with classical invariants $(\operatorname{tb},\operatorname{rot})=(-6,-1)$ in Figure 3. The fundamental 4-Legendrian rack is the quotient of the free 4-Legendrian rack $\langle x_{1},x_{2},x_{3}\rangle$ by the congruence relation generated by the relations
@@
A front projection of a Legendrian left handed trefoil. The arcs are labeled in a way that corresponds to the fundamental 4-Legendrian rack.
Figure 3: Legendrian left handed trefoil with arcs labeled.
6 Main Theorem
@@
This idea is similar to an approach Cheng and He took to prove the main theorem of [2]. Namely, they observed that every relation in the canonical presentation of the fundamental GL-rack of $L$ can be written in the form
@@
$x=\underbrace{(d_{l}d_{r}\dots d_{l}d_{r})}_{\operatorname{rot}(L)\text{ times}}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x)\hskip 18.06749pt\text{or}\hskip 18.06749ptx=\underbrace{(d_{r}d_{l}\dots d_{r}d_{l})}_{\operatorname{rot}(L)\text{ times}}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x).$
@@
Since $L$ is a Legendrian knot, $U+D$ must be even, and $U,D>0$. It follows from Proposition 6.1 that one of $u_{l}d_{r}$, $d_{r}u_{l}$, $u_{r}d_{l}$, or $d_{l}u_{r}$ appears in $W$. Introduce a $\sigma\sigma^{-1}$ in the middle of this pair. Since $\sigma$ must commute with $u_{l},u_{r},d_{l},\text{ and }d_{r}$, we can commute the $\sigma^{-1}$ to the rightmost part of the word. By Example 4.9, we can then use the $\sigma$ to cancel out the $u_{l}d_{r}$, $d_{r}u_{l}$, $u_{r}d_{l}$, or $d_{l}u_{r}$. We can repeat this action until there are no longer any upward cusp functions in our word.
The word now only contains $D-U=2\operatorname{rot}(L)$ total $d_{l}$’s and $d_{r}$’s. Either $d_{l}$ or $d_{r}$ begin the word and the two functions must alternate. Meaning we are left with $\operatorname{rot}(L)$ pairs of $d_{l}d_{r}$’s or $d_{r}d_{l}$’s followed by $w(L)-U=\operatorname{rot}(L)+\operatorname{tb}(L)$ total $\sigma$’s in our word, as desired. ∎
@@
44-Legendrian permutation racks cannot distinguish isotopy classes of Legendrian knots with the same classical invariants.
@@
Let $(X_{\sigma},u_{l},u_{r},d_{l},d_{r})$ be an any 44-Legendrian permutation rack. Assume $L_{1}$ and $L_{2}$ are two distinct Legendrian isotopy classes of knots with the same classical invariants. Fix oriented front projections of $L_{1}$ and $L_{2}$.
@@
$x=\underbrace{(d_{l}d_{r}\dots d_{l}d_{r})}_{\operatorname{rot}(L)\text{ times}}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x)\hskip 18.06749pt\text{or}\hskip 18.06749ptx=\underbrace{(d_{r}d_{l}\dots d_{r}d_{l})}_{\operatorname{rot}(L)\text{ times}}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x).$
We can choose x in each respective front projection such that the word made of $d_{l}$’s and $d_{r}$’s begins with the same function. Since $L_{1}$ and $L_{2}$ share the same classical invariants, it follows that these two relations are identical. Hence, the number of homomorphisms from the fundamental 44-Legendrian rack to $(X_{\sigma},u_{l},u_{r},d_{l},d_{r})$ will be equal for both $L_{1}$ and $L_{2}$. ∎

### v2

@@ 4.1 Legendrian Knots
Legendrian knots are important objects of study in contact topology. We briefly review several concepts from the theory. For a formal treatment, we refer the reader to the survey of Etnyre [Etn05].
The standard contact structure is the kernel of the differential 1-form dzydxdz-y\,dx in $\mathbb{R}^{3}$, which is depicted as a plane field in Figure 1. A smooth knot in $\mathbb{R}^{3}$ is called Legendrian if it lies everywhere tangent to the standard contact structure.
An assignment of a plane to each point in Euclidean three-space. The planes are horizontal along the x-axis and twist in the y-direction. They approach verticality, but never reach it.
@@
Legendrian knots are usually studied via their projections to the $xz$-plane viewed from the negative $y$-direction. These are called front projections. Front projections of Legendrian knots have two key features that distinguish them from projections of smooth knots. First, since tangent lines can never be vertical, front projections have cusps in place of vertical tangencies. Second, due to the direction of twisting, the overstrand at each crossing is always the strand having the more negative slope. Note that an oriented front projection never has two leftward-oriented or rightward-oriented cusps placed adjacent to each other. For example, Figure 2 (cf. Figure 4) depicts a Legendrian left-handed trefoil in its front projection.
An oriented front projection of an Legendrian left-handed trefoil.
Figure 2: Oriented front projection of a Legendrian left-handed trefoil.
@@
However, there exist many nonequivalent Legendrian knots that share the same topological knot type and classical invariants; see, for example, [Etn05, Sec. 4]. This motivates the search for Legendrian knot invariants that distinguish such pairs of Legendrian knots.
@@
In particular, Kimura [Kim23] and Cheng and He [CH25] posed the open questions of whether GL-racks and, more generally, 4-Legendrian racks provide effective invariants of Legendrian knots. While this question remains open in general, Kimura [Kim24, Thm. 4.2.3] gave a negative answer for all 4-Legendrian quandles, and our main theorem (Theorem 5.5) provides a negative answer for all 4-Legendrian permutation racks.
@@ 4.2 Coloring Invariants
In 2024, Kimura [Kim24, Sec. 4.2] introduced an invariant of Legendrian knots $L$ called the fundamental 4-Legendrian rack, which we denote by $F(L)$. This is defined in the same vein as the fundamental GL-rack of $L$, which Karmakar, Saraf, and Singh introduced in [KSSar, Thm. 4.3], and the fundamental quandle of a smooth knot, which was independently introduced by Joyce [Joy82] and Matveev [Mat84].
We outline the construction here; see [Kim24, Sec. 4.2] for details, and cf. [KSSar, CH25]. View 4-Legendrian racks as an algebraic theory with two binary operations $\triangleright^{\pm 1}$ and four unary operations $u_{l}^{\pm 1},u_{r}^{\pm 1}$. By general results from universal algebra, free 4-Legendrian racks exist and are characterized up to isomorphism by the usual universal property. Moreover, we can take quotients of 4-Legendrian racks by congruence relations.
Fix an oriented front projection of a Legendrian knot $L$. Label the arcs (i.e., connected components) of the front projection by $x_{1},\dots,x_{n}$, and let $F\coloneq\langle x_{1},\dots,x_{n}\rangle$ be the free 4-Legendrian rack generated by $x_{1},\dots,x_{n}$. Starting at any crossing in the front projection, traverse the knot using its given orientation. At each cusp, impose a label $u_{l},u_{r},d_{l}$, or $d_{r}$ depending on the orientation of the cusp in the obvious way; cf. (3.1). At each crossing, impose a relation on $F$ between arcs as illustrated in Figure 3. Define $F(L)$ to be the quotient of $F$ by the congruence relation generated by these $n$ crossing relations.
@@
One can show that the isomorphism class of $F(L)$ is an invariant of $L$ by using Legendrian Reidemeister moves; see [Kim24, Sec. 4.2]. Of course, it is difficult to distinguish infinite 4-Legendrian racks directly from their generators and relations, so we instead consider colorings of $F(L)$ by finite 4-Legendrian racks. These are easier to compute since $F(L)$ is finitely presented (having $n$ generators and $n$ relations in the above construction).
Definition 4.4 ([Kim24, Def. 23]).
@@
Since $F(L)$ is an invariant of $L$, the set $\operatorname{Hom}(F(L),X)$ of colorings of $L$ by a given 4-Legendrian rack $(X,\triangleright,u_{l},u_{r})$ is also an invariant of $L$; cf. [Kim24, Rem. 30].
@@
Let $L$ be the Legendrian left-handed trefoil with classical invariants $(\operatorname{tb},\operatorname{rot})=(-6,-1)$ in Figure 2. By following the procedure described above, we obtain Figure 4. Thus, the fundamental 4-Legendrian rack $F(L)$ is the quotient of the free 4-Legendrian rack $\langle x_{1},x_{2},x_{3}\rangle$ by the congruence relation generated by the relations
@@
An oriented front projection of an Legendrian left-handed trefoil. The arcs are labeled in a way that corresponds to the fundamental 4-Legendrian rack.
Figure 4: Legendrian left-handed trefoil with arcs labeled.
5 Proof of the Main Theorem
@@
This idea is similar to an approach Cheng and He took to prove the main theorem of [CH25]. Namely, they observed that every relation in the canonical presentation of the fundamental GL-rack of $L$ can be written in the form
@@
$x=(d_{l}d_{r})^{\operatorname{rot}(L)}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x)\qquad\text{or}\qquad x=(d_{r}d_{l})^{\operatorname{rot}(L)}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x).$
@@
Since $L$ is a Legendrian knot, $U+D$ must be even, and $U,D>0$. It follows from Proposition 5.1 that one of $u_{l}d_{r}$, $d_{r}u_{l}$, $u_{r}d_{l}$, or $d_{l}u_{r}$ appears in $W$. Introduce a $\sigma\sigma^{-1}$ in the middle of this pair. Since $\sigma$ must commute with $u_{l},u_{r},d_{l},\text{ and }d_{r}$, we can commute the $\sigma^{-1}$ to the rightmost part of the word. By Example 3.7, we can then use the $\sigma$ to cancel out the $u_{l}d_{r}$, $d_{r}u_{l}$, $u_{r}d_{l}$, or $d_{l}u_{r}$. We can repeat this action until there are no longer any upward cusp functions $u_{l},u_{r}$ in our word.
The word now only contains $D-U=2\operatorname{rot}(L)$ total $d_{l}$’s and $d_{r}$’s. Either $d_{l}$ or $d_{r}$ begin the word, and these two functions must alternate. This leaves us with $\operatorname{rot}(L)$ pairs of $d_{l}d_{r}$’s or $d_{r}d_{l}$’s followed by $w(L)-U=\operatorname{rot}(L)+\operatorname{tb}(L)$ total $\sigma$’s in our word, as desired. ∎
As an aside, Lemma 5.2 allows us to read $\operatorname{rot}(L)$ and $\operatorname{tb}(L)$ off of the relations of a coloring $F(L)\to X_{\sigma}$. This yields the following, which may be seen as a stronger version of the main theorem of [CH25] for 4-Legendrian racks.
Corollary 5.3 (Theorem 1.2).
The fundamental 4-Legendrian rack $F(L)$ recovers the classical invariants of $L$.
Remark 5.4.
Some authors consider the topological knot type to be a classical invariant of Legendrian knots. The statement of Corollary 5.3 still holds under this convention. To see this, recall that the fundamental quandle is a complete invariant of smooth knots (see [Mat84, Joy82]). Given a Legendrian embedding $L$ of a knot $K$, we can recover the fundamental quandle of $K$ from $F(L)$ by imposing the congruence relation generated by $x\sim u_{l}(x)\sim d_{r}(x)\sim u_{r}(x)\sim d_{l}(x)$ for all $x\in F(L)$; cf. [KSSar, Rem. 4.6].
@@
Let $(X_{\sigma},u_{l},u_{r},d_{l},d_{r})$ be any 4-Legendrian permutation rack. Assume $L_{1}$ and $L_{2}$ are two distinct Legendrian isotopy classes of knots with the same classical invariants. Fix oriented front projections of $L_{1}$ and $L_{2}$.
@@
$x=(d_{l}d_{r})^{\operatorname{rot}(L)}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x)\qquad\text{or}\qquad x=(d_{r}d_{l})^{\operatorname{rot}(L)}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x).$
We can choose $x$ in each respective front projection such that the word made of $d_{l}$’s and $d_{r}$’s begins with the same function. Since $L_{1}$ and $L_{2}$ share the same classical invariants, it follows that these two relations are identical. Hence, by the definition of a 4-Legendrian rack coloring, the sets of colorings of $L_{1}$ and $L_{2}$ by $(X_{\sigma},u_{l},u_{r},d_{l},d_{r})$ are equal. ∎
6 Areas for Further Inquiry
We conclude by proposing areas for further research. These complement the open question of whether there exist 4-Legendrian racks that can distinguish Legendrian knots with the same classical invariants.
  • Study the coloring invariant of 4-Legendrian racks that are not permutation racks.
  • Extend the algebraic results for GL-racks in [Ta25a] to 4-Legendrian racks.
  • Develop enhancements of the 4-Legendrian rack coloring invariant.
  • Study the algebraic properties and distinguishing power of 4-Legendrian biracks, which Kimura introduced in his PhD thesis [Kim24].
  • Develop rack-theoretic invariants of Legendrian graphs.
  • Develop rack-theoretic invariants of transverse knots.